OpenStudy (anonymous):

convert to fraction in lowest terms 42.8%

Parth (parthkohli):

=> 428/1000 => 214/500 => 107/250

OpenStudy (anonymous):

% = /100 so 42.8/100 You can multiply both by 10 to get rid of that decimal point so 428/1000 Then find common terms you can use to simplify it.

Parth (parthkohli):

107/250 is the simplest form. DO you get it?

OpenStudy (anonymous):

yes got thank you

OpenStudy (anonymous):

42.8%=42.8*(1/100)=428/1000=214/500=107/250 this is the lowest fraction